# Bayesian Characterization and Identifiability of Multilevel Quantum Systems

### v1

@@ 4.2 Likelihood Function
For clarity of presentation, we denote by $\widetilde{\mathrm{\boldsymbol{\theta}}}$ the true distribution of the parameters (i.e. the parameters of the quantum testbed). For a total of $N$ independent shots, let $y_{i}\sim\mathrm{Pr}{}(\mathrm{\boldsymbol{y}}|\widetilde{\mathrm{\boldsymbol{\theta}}};\mathrm{\boldsymbol{\xi}})$ be an outcome of a shot (i.e., the state measured for shot $i$) after applying control $\mathrm{\boldsymbol{\xi}}$, $n_{i}$ the number of times $y_{i}$ is observed with $\sum_{i}n_{i}=N$. Then, the likelihood can be modeled by a multinomial distribution [17, 35] defined by
@@
where $\begin{pmatrix}N\\ n_{1},n_{2},\ldots,n_{M}\end{pmatrix}:=\dfrac{N!}{\prod_{i=1}^{\left|{\mathbb{S}}\right|}n_{i}!}$ is the multinomial coefficient and $\mathrm{Pr}{}(y_{i}|\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}})=\mathrm{Pr}{}(y_{i}|\rho(\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}}))\approx\dfrac{n_{i}}{N}$ is the probability of observing $y_{i}$ from Eq. (7). It should be mentioned that the multinomial coefficient can be omitted since the weights in Eq. (24) are renormalized at each epoch. For a large number of draws $N$, this omission may lead to numerical underflow as the resulting expression would produce near zero probabilities. Alternatively, the multinomial distribution asymptotically tends to a Gaussian distribution [37]
@@
where $\Sigma$ is the covariance that accounts for errors between the two models. This model of the likelihood has been used extensively in Bayesian experimental design [34, 38, 35] and in Markov Chain Monte Carlo (MCMC) methods.
@@ 5 Case Studies for Quantum Characterization
We now present the problem formulation for verification of our characterization technique using synthetic test problems. We consider an open system modeled by the Lindblad master equation with true value of the parameter set $\mathrm{\boldsymbol{\theta}}=\left\{\omega,\chi,T_{1},T_{2}\right\}$ to be that of the standard QPU in Quantum Design and Integration Testbed (QuDIT) at LLNL [39]. We mainly consider uncertainties in these four parameters, which are the most pronounced noise sources in the experiments, especially for a quantum gate in multi-level system. The QPU has a single transmon in 3D aluminum cavity whose parameters are given below.
@@
In addition to the noise introduced by imposing a distribution over the model parameters, we also introduce shot noise by sampling a finite number of outcomes $N$ (in Eq. (29)). When omitting shot noise and using exact $\mathrm{Pr}{}(\mathrm{\boldsymbol{y}}|\widetilde{\mathrm{\boldsymbol{\theta}}};\mathrm{\boldsymbol{\xi}})$ (i.e. when $N\rightarrow\infty$), we employ the Gaussian model for the likelihood (Eq. (30)). We refer to those experiments performed using point-estimates of parameters and exact measurements (i.e. $N=\infty$) as noise-free experiments. We use a total of 2000 Monte Carlo particles and limit the total number of epochs to 500. It was previously mentioned that the topology of the utility function can be highly multimodal, hence we employ the non-gradient based differential evolution algorithm, since it is easily parallelizable and less susceptable to convergence to local extrema. Each experiment will be defined by a different control pulse that is parameterized as a sequence of piecewise constant segments in the rotating frame, with a drive frequency $\omega_{d}$ and each segment defined by $p(t)$, $q(t)$ and $\Delta t$ as shown in Fig. 2. Hence, for an $N_{p}$ segment pulse, the total number of parameters defining the experiment are $N_{T}=3N_{p}+1$. The bounds on the pulse parameters are $p(t),q(t)\in[-12,12]$ MHz, $\Delta t\in[1,30]~{}\mu$s and $\omega_{d}\in[3.5,4.5]$. We assume each experiment begins with the qutrit in the ground state $\left|0\right\rangle$.
@@
6 Results: Characterizing Multilevel Quantum System
We consider Bayesian experimental design under a fixed control pulse and an adaptive control pulse parameterization. In the fixed case, the number of control pulse segments remains constant, whereas in the adaptive case, the number of pulse segments vary with epochs.
@@ 6.1 Fixed Pulse Parameterization
We first investigate the effectiveness of our approach under a fixed pulse parameterization and the problem formulation presented in Sec. 5. Figure 4 shows the mean error and the standard deviation of the estimated distributions using different number of control pulse segments under different sources of noise. Immediately we see that the mean error and the standard deviation for $\omega$ and $\chi$ distributions are orders of magnitude smaller than those for the distributions of $T_{1}$ and $T_{2}$. Furthermore, variances of all four parameter distributions decrease near-monotonically. In both the noise and noise-free case, the transition frequency and anharmonicity show a faster rate of convergence than the decoherence parameters, with several orders of magnitude reduction in only the first few epochs. We see a better overall convergence in $T_{1}$ and $T_{2}$ when using a more robust pulse parameterization with a larger number of segments, but with a slight decrease in the accuracy of $\omega$ and $\chi$ estimates. Control pulses with a greater number of segments lead to longer pulse durations and better expose the dissipative dynamics that are present on larger timescales. This suggests that an adaptive pulse parameterization - where short pulse sequences are used to accurately estimate $\omega$ and $\chi$, then longer sequences are used to improve the $T_{1}$ and $T_{2}$ estimates - may be beneficial.
@@
Figure 4: Convergence of mean error $\delta(\widehat{\mu})$ and standard deviation $\widehat{\sigma}$ of the model parameters when performing noise-free (top) experiments and experiments with parameter noise (bottom) using pulses with: (a,c) three segments and (b,d) six segments.
@@
Figure 5 shows the exact (imposed) and the converged distribution for the four parameters using a different number of pulse segments $N_{p}$ and shots $N_{s}$. We see a good agreement in the distributions for $\omega$ and $\chi$ whereas the distributions of $T_{1}$ and $T_{2}$ feature larger variances. The mean of $T_{1}$ has better agreement with that of the exact distribution, whereas the distribution of $T_{2}$ displays both a larger error in the mean and a wider spread. As expected, the noise introduced by sampling a finite number of shots introduces additional errors; this often manifests as a larger shift in mean or larger variances. Furthermore, we see that shorter pulse sequences yield better estimates for $\omega$ and $\chi$ whereas longer pulse sequences result in better estimates for $T_{1}$ and $T_{2}$. The statistical moments of the converged distribution reported in Tab. 2 show that the frequency parameters are less susceptable to shot noise than the decoherence times.
@@
Figure 5: Converged parameter distribution when performing noisy experiments using a different number of pulse segments ($N_{p}$) and different number of shots $N_{s}$. Note the distributions are translated by the mean of the exact distributions.
@@
Noise Type Segments $\omega$ (GHz) $\chi$ (MHz) $T_{1}$ ($\mu$s) $T_{2}$ ($\mu$s)
None 3 4.0108 (8.950E-07) 127.8 (3.672E-06) 43.806 (2.397) 26.011 (1.789)
6 4.0108 (1.192E-06) 127.8 (5.458E-06) 45.534 (2.255) 23.936 (1.472)
Parameter 3 4.0108 (7.169E-07) 127.8 (2.100E-06) 45.289 (2.387) 26.052 (1.624)
6 4.0108 (1.801E-06) 127.8 (7.796E-06) 44.677 (3) 23.308 (1.142)
Parameter and Shot 3 4.0108 (7.065E-07) 127.8 (2.335E-06) 46.738 (3.081) 25.123 (2.169)
6 4.0108 (1.234E-06) 127.8 (4.794E-06) 45.584 (1.805) 22.063 (0.8545)
When model parameters are not unique, quantifying the accuracy using only errors in the parameters is not sufficient. In such cases, it is imperative to quantify the accuracy of model response/predictions. Figure 6 shows the probability density of the errors in the population due to different noise types, pulse segments, and number of shots. Furthermore, the mean and standard deviation of the errors is also presented in Tab. 3. In the case of noise-free experiments, we see that longer pulse sequences yield model parameters that result in more accurate predictions (i.e. higher densities for smaller errors). This is due to the more accurate estimation of $T_{1}$ and $T_{2}$ and therefore, more accurate representation of the dissipative dynamics that are present in longer duration pulses (i.e., larger gate repitition). When performing estimation using noisy experiments, we see that the error distribution for all population and pulse parameterization are similar, and are an order of magnitude larger than that of noise-free predictions. Furthermore, despite the lack of accuracy in the decoherence parameters compared to the frequency parameters, the forward model with estimated parameters yields sufficiently accurate predictions with mean errors on the order of $10^{-4}$ and variances on the order of $10^{-3}$.
Figure 6: Distribution of the error in the population when using a different number of pulse segments and performing: a) noise-free experiments, b) experiments with parameter noise, and c) experiments with parameter and shot noise.
@@
Noise Type Segments $\left|0\right\rangle$ $\left|1\right\rangle$ $\left|2\right\rangle$
None 3 5.31E-05 (1.60E-03) -6.39E-05 (1.20E-03) 1.08E-05 (1.53E-03)
6 8.60E-06 (5.93E-04) -1.98E-05 (4.08E-04) 1.12E-05 (5.47E-04)
Parameter 3 1.18E-04 (2.56E-03) -1.87E-04 (2.10E-03) 6.85E-05 (2.49E-03)
6 -6.36E-05 (3.03E-03) 1.09E-04 (2.21E-03) -4.58E-05 (2.78E-03)
Parameter and Shot 3 1.17E-04 (2.15E-03) -2.03E-04 (1.88E-03) 8.58E-05 (2.11E-03)
6 -1.07E-04 (2.99E-03) 1.61E-04 (2.31E-03) -5.40E-05 (2.86E-03)
6.2 Adaptive Pulse Parameterization
Since different pulse parameterization greatly affects the accuracy of the estimated parmeter distributions asymmetrically, we now investigate an adaptive pulse parameterization to improve convergence in all four parameters. We begin with three segments and increase the number of segments by one every 100 epochs. We only consider cases with parameter (defined as random variables) and shot noise, with 500, 1000, and $\infty$ shots. Figure 7 shows the convergence history in the mean and the standard deviation of the four parameters. Although the accuracy of the frequencies is comparable to the fixed pulse parameterization, the decoherence parameters are better estimated with the adaptive pulse parameterization. The mean error in the decoherence parameters are an order of magnitude lower than the fixed pulse parameterization. When we introduce shot noise, we see the accuracy is comparable to the fixed pulse parameterization. Table 4 shows the mean and standard deviation of the four parameters. The improved accuracy of $T_{1}$ and $T_{2}$ estimation is clear using a moderate number of shots. As expected, large sampling errors (from severe undersampling) can greatly affect the accuracy of the parameter estimates.
Figure 7: Convergence of mean error $\delta(\widehat{\mu})$ and standard deviation $\widehat{\sigma}$ of the model parameters when performing experiments with parameter noise using : (a) 500 shots, (b) 1000 shots and (c) $\infty$ shots.
Table 4: Mean and standard deviation (in the brackets) of the parameters using an adaptive pulse parameterization

### v2

@@ 4.2 Likelihood Function
For clarity of presentation, we denote by $\widetilde{\mathrm{\boldsymbol{\theta}}}$ the true distribution of the parameters (i.e. the parameters of the quantum testbed). For a total of $N$ independent shots, let $y_{i}\sim\mathrm{Pr}{}(\mathrm{\boldsymbol{y}}|\widetilde{\mathrm{\boldsymbol{\theta}}};\mathrm{\boldsymbol{\xi}})$ be a possible outcome of a shot (i.e., the state measured for a shot) after applying control $\mathrm{\boldsymbol{\xi}}$, $n_{i}$ the number of times $y_{i}$ is observed with $\sum_{i}n_{i}=N$. Then, the likelihood can be modeled by a multinomial distribution [17, 38] defined by
@@
where $\begin{pmatrix}N\\ n_{1},n_{2},\ldots,n_{M}\end{pmatrix}:=\dfrac{N!}{\prod_{i=1}^{\left|{\mathbb{S}}\right|}n_{i}!}$ is the multinomial coefficient, $y_{i}\sim\mathrm{Pr}{}(y_{i}|\widetilde{\mathrm{\boldsymbol{\theta}}};\mathrm{\boldsymbol{\xi}})=\mathrm{Pr}{}(y_{i}|\rho(\widetilde{\mathrm{\boldsymbol{\theta}}};\mathrm{\boldsymbol{\xi}}))\approx\dfrac{n_{i}}{N}$ is the probability, estimated from experiments, of observing $y_{i}$ and $\mathrm{Pr}{}(y_{i}|\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}})$ is the probability of observing $y_{i}$ computed using Eq. (7) and the numerical model for a given $\mathrm{\boldsymbol{\theta}}$ and $\mathrm{\boldsymbol{\xi}}$. It should be mentioned that the multinomial coefficient can be omitted since the weights in Eq. (24) are renormalized at each epoch. For a large number of draws $N$, this omission may lead to numerical underflow as the resulting expression would produce near zero probabilities, particularly when $\mathrm{Pr}{}(y_{i}|\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}})$ is small. Alternatively, under certain assumptions, the multinomial distribution asymptotically tends to a Gaussian distribution [42]
@@
where $\Sigma$ is the covariance that accounts for errors between the numerical model and the ‘true’ model (i.e. experiment). This model of the likelihood has been used extensively in Bayesian experimental design [37, 43, 38] and in Markov Chain Monte Carlo (MCMC) methods. The exact form of $\Sigma$ is not known a priori and must be estimated. Let $\mathrm{\boldsymbol{\mathcal{Y}}}(\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}})=\left\{\mathrm{Pr}{}(y_{i}|\mathrm{\boldsymbol{\theta}};\mathrm{\boldsymbol{\xi}}),i=1,\ldots,N\right\}$ and $\widetilde{\mathrm{\boldsymbol{\mathcal{Y}}}}(\mathrm{\boldsymbol{\xi}})=\left\{\mathrm{Pr}{}(y_{i}|\widetilde{\mathrm{\boldsymbol{\theta}}};\mathrm{\boldsymbol{\xi}}),i=1,\ldots,N\right\}$ be vectors of probabilities computed using the numerical model and experimentally measured, respectively, after applying control $\mathrm{\boldsymbol{\xi}}$. Then, we estimate $\Sigma$ by means of Monte Carlo sampling over the prior for $\mathrm{\boldsymbol{\theta}}$ with $N_{\mathrm{\boldsymbol{\theta}}}$ samples and random control $\mathrm{\boldsymbol{\xi}}$ with $N_{\mathrm{\boldsymbol{\xi}}}$ samples. The sample covariance matrix is then computed as
$\Sigma=\dfrac{1}{N_{\mathrm{\boldsymbol{\theta}}}}\dfrac{1}{N_{\mathrm{\boldsymbol{\xi}}}}\sum_{i=1}^{N_{\mathrm{\boldsymbol{\theta}}}}\sum_{j=1}^{N_{\mathrm{\boldsymbol{\xi}}}}\left(\mathrm{\boldsymbol{\mathcal{Y}}}(\mathrm{\boldsymbol{\theta}}_{i};\mathrm{\boldsymbol{\xi}}_{j})-\widetilde{\mathrm{\boldsymbol{\mathcal{Y}}}}(\mathrm{\boldsymbol{\xi}}_{j})\right)\left(\mathrm{\boldsymbol{\mathcal{Y}}}(\mathrm{\boldsymbol{\theta}}_{i};\mathrm{\boldsymbol{\xi}}_{j})-\widetilde{\mathrm{\boldsymbol{\mathcal{Y}}}}(\mathrm{\boldsymbol{\xi}}_{j})\right)^{T}$ (31)
@@ 5 Case Studies for Quantum Characterization
We now present the problem formulation for verification of our characterization technique using synthetic test problems. We consider an open system modeled by the Lindblad master equation with true value of the parameter set $\mathrm{\boldsymbol{\theta}}=\left\{\omega,\chi,T_{1},T_{2}\right\}$ to be that of the standard QPU in Quantum Design and Integration Testbed (QuDIT) at LLNL [44]. We mainly consider uncertainties in these four parameters, which are the most pronounced noise sources in the experiments, especially for a quantum gate in multi-level system. The QPU has a single transmon in 3D aluminum cavity whose parameters are given in Tab. 1.
@@
Since the ‘true’ distribution is not known in practice, the use of an uninformative prior, like the uniform distribution, avoids bias in the inference. In addition to the uncertainty introduced by imposing a distribution over the model parameters, we also introduce shot noise (i.e. sampling uncertainty) by sampling a finite number of outcomes $N$ (in Eq. (29)). When omitting shot noise and using exact $\mathrm{Pr}{}(\mathrm{\boldsymbol{y}}|\widetilde{\mathrm{\boldsymbol{\theta}}};\mathrm{\boldsymbol{\xi}})$ (i.e. when $N\rightarrow\infty$), we employ the Gaussian model for the likelihood (Eq. (30)). We refer to those experiments performed using point-estimates of parameters and exact measurements (i.e. $N=\infty$) as exact experiments and those experiments with uncertainty (either parameter or sampling) as stochastic experiments. We use a total of 2000 Monte Carlo particles and limit the total number of epochs to 500. The results with different number of particles is given in Appendix A. It was previously mentioned that the topology of the utility function can be highly multimodal [39], hence we employ the non-gradient based differential evolution algorithm, since it is easily parallelizable and less susceptable to convergence to local extrema. Each experiment will be defined by a different control pulse that is parameterized as a sequence of piecewise constant segments in the rotating frame, with a drive frequency $\omega_{d}$ and each segment defined by $p(t)$, $q(t)$ and $\Delta t$ as shown in Fig. 2. Although other pulse parameterization, such as spline parameterization [45], can also be used, the piecewise-constant pulses simplies matrix exponentiation since it can be performed for each segment simultaneously and in parallel. For a finely resolved pulse, this piecewise-constant parameterization will lead to an overwhelming number of parameters that must be optimized, and therefore, a spline-based parameterization is more applicable. Hence, for an $N_{p}$ segment pulse, the total number of parameters defining the experiment are $N_{T}=3N_{p}+1$. The bounds on the pulse parameters are $p(t),q(t)\in[-12,12]$ MHz, $\Delta t\in[1,30]~{}\mu$s and $\omega_{d}\in[3.5,4.5]$. We assume each experiment begins with the qutrit in the ground state $\left|0\right\rangle$.
@@
6 Results: Characterizing Qutrit Systems
We consider Bayesian experimental design under a fixed control pulse and an approach for iterative control pulse extension. In the fixed case, the number of control pulse segments remains constant, whereas in the iterative pulse extension case, the number of pulse segments increase with epochs.
@@ 6.1 Fixed Pulse Parameterization
We first investigate the effectiveness of our approach under a fixed pulse parameterization and the problem formulation presented in Sec. 5. Figure 4 shows the mean error and the standard deviation of the estimated distributions using different number of control pulse segments under different sources of uncertainty. Immediately we see that the mean error and the standard deviation for $\omega$ and $\chi$ distributions are orders of magnitude smaller than those for the distributions of $T_{1}$ and $T_{2}$. Furthermore, variances of all four parameter distributions decrease near-monotonically. In both the exact and stochastic experiments, the transition frequency and anharmonicity show a faster rate of convergence than the decoherence parameters, with several orders of magnitude reduction in only the first few epochs. We see a better overall convergence in $T_{1}$ and $T_{2}$ when using a more robust pulse parameterization with a larger number of segments, but with a slight decrease in the accuracy of $\omega$ and $\chi$ estimates. Control pulses with a greater number of segments lead to longer pulse durations and better expose the dissipative dynamics that are present on larger timescales. This suggests that an adaptive pulse parameterization - where short pulse sequences are used to accurately estimate $\omega$ and $\chi$, then longer sequences are used to improve the $T_{1}$ and $T_{2}$ estimates - may be beneficial.
@@
Figure 4: Convergence of mean error $\delta(\widehat{\mu})$ and standard deviation $\widehat{\sigma}$ of the model parameters when performing exact (top) experiments and experiments with parameter uncertainty (bottom) using pulses with: (a,c) three segments and (b,d) six segments. The solid and dashed lines represent the convergence in mean error and standard deviation, respectively.
@@
Figure 5 shows the exact (imposed) and the converged distribution for the four parameters using a different number of pulse segments $N_{p}$ and shots $N_{s}$. We see a good agreement in the distributions for $\omega$ and $\chi$ whereas the distributions of $T_{1}$ and $T_{2}$ feature larger variances. The mean of $T_{1}$ has better agreement with that of the exact distribution, whereas the distribution of $T_{2}$ displays both a larger error in the mean and a wider spread. As expected, the uncertainty introduced by sampling a finite number of shots introduces additional errors; this often manifests as a larger shift in mean or larger variances. Furthermore, we see that shorter pulse sequences yield better estimates for $\omega$ and $\chi$ whereas longer pulse sequences result in better estimates for $T_{1}$ and $T_{2}$. The statistical moments of the converged distribution reported in Tab. 2 show that the frequency parameters are less susceptable to shot noise than the decoherence times.
@@
Figure 5: Converged parameter distribution when performing stochastic experiments using a different number of pulse segments ($N_{p}$) and different number of shots $N_{s}$. Note the distributions are translated by the mean of the exact distributions.
@@
Uncertainty Type Segments $\omega$ (GHz) $\chi$ (MHz) $T_{1}$ ($\mu$s) $T_{2}$ ($\mu$s)
None 3 4.0108 (8.950×$\times 10^{-7}$) 127.8 (3.672×$\times 10^{-6}$) 43.806 (2.397) 26.011 (1.789)
6 4.0108 (1.192×$\times 10^{-6}$) 127.8 (5.458×$\times 10^{-6}$) 45.534 (2.255) 23.936 (1.472)
Parameter 3 4.0108 (7.169×$\times 10^{-7}$) 127.8 (2.100×$\times 10^{-6}$) 45.289 (2.387) 26.052 (1.624)
6 4.0108 (1.801×$\times 10^{-6}$) 127.8 (7.796×$\times 10^{-6}$) 44.677 (3) 23.308 (1.142)
Parameter and Shot 3 4.0108 (7.065×$\times 10^{-7}$) 127.8 (2.335×$\times 10^{-6}$) 46.738 (3.081) 25.123 (2.169)
6 4.0108 (1.234×$\times 10^{-6}$) 127.8 (4.794×$\times 10^{-6}$) 45.584 (1.805) 22.063 (0.8545)
When model parameters are not unique, quantifying the accuracy using only errors in the parameters is not sufficient. In such cases, it is imperative to quantify the accuracy of model response/predictions. Figure 6 shows the probability density of the errors in the population due to different sources of uncertainty, pulse segments, and number of shots. Furthermore, the mean and standard deviation of the errors is also presented in Tab. 3. In the case of exact experiments, we see that longer pulse sequences yield model parameters that result in more accurate predictions (i.e. higher densities for smaller errors). This is due to the more accurate estimation of $T_{1}$ and $T_{2}$ and therefore, more accurate representation of the dissipative dynamics that are present in longer duration pulses (i.e., larger gate repitition). When performing estimation using stochastic experiments, we see that the error distribution for all population and pulse parameterization are similar, and are an order of magnitude larger than that of exact predictions. Furthermore, despite the lack of accuracy in the decoherence parameters compared to the frequency parameters, the mathematical model with estimated parameters yields sufficiently accurate predictions with mean errors on the order of $10^{-4}$ and variances on the order of $10^{-3}$.
Figure 6: Distribution of the error in the population when using a different number of pulse segments and performing: a) exact experiments, b) experiments with parameter uncertainty, and c) experiments with parameter uncertainty and shot noise.
@@
Uncertainty Type Segments $\left|0\right\rangle$ $\left|1\right\rangle$ $\left|2\right\rangle$
None 3 5.31×$\times 10^{-5}$ (1.60×$\times 10^{-3}$) -6.39×$\times 10^{-5}$ (1.20×$\times 10^{-3}$) 1.08×$\times 10^{-5}$ (1.53×$\times 10^{-3}$)
6 8.60×$\times 10^{-6}$ (5.93×$\times 10^{-4}$) -1.98×$\times 10^{-5}$ (4.08×$\times 10^{-4}$) 1.12×$\times 10^{-5}$ (5.47×$\times 10^{-4}$)
Parameter 3 1.18×$\times 10^{-4}$ (2.56×$\times 10^{-3}$) -1.87×$\times 10^{-4}$ (2.10×$\times 10^{-3}$) 6.85×$\times 10^{-5}$ (2.49×$\times 10^{-3}$)
6 -6.36×$\times 10^{-5}$ (3.03×$\times 10^{-3}$) 1.09×$\times 10^{-4}$ (2.21×$\times 10^{-3}$) -4.58×$\times 10^{-5}$ (2.78×$\times 10^{-3}$)
Parameter and Shot 3 1.17×$\times 10^{-4}$ (2.15×$\times 10^{-3}$) -2.03×$\times 10^{-4}$ (1.88×$\times 10^{-3}$) 8.58×$\times 10^{-5}$ (2.11×$\times 10^{-3}$)
6 -1.07×$\times 10^{-4}$ (2.99×$\times 10^{-3}$) 1.61×$\times 10^{-4}$ (2.31×$\times 10^{-3}$) -5.40×$\times 10^{-5}$ (2.86×$\times 10^{-3}$)
6.2 Iterative Pulse Extension
Since different pulse parameterization greatly affects the accuracy of the estimated parmeter distributions asymmetrically, we now investigate iterative pulse extensions to improve convergence in all four parameters. We begin with three segments and increase the number of segments by one every 100 epochs. We only consider cases with parameter (defined as random variables) and shot noise, with 500, 1000, and $\infty$ shots. Figure 7 shows the convergence history in the mean and the standard deviation of the four parameters. Although the accuracy of the frequencies is comparable to the fixed pulse parameterization, the decoherence parameters are better estimated with iterative pulse extensions. The mean error in the decoherence parameters are an order of magnitude lower than the fixed pulse parameterization. When we introduce shot noise, we see the accuracy is comparable to the fixed pulse parameterization. Table 4 shows the mean and standard deviation of the four parameters. The improved accuracy of $T_{1}$ and $T_{2}$ estimation is clear using a moderate number of shots. As expected, large sampling errors (from severe undersampling) can greatly affect the accuracy of the parameter estimates.
Figure 7: Convergence of mean error $\delta(\widehat{\mu})$ and standard deviation $\widehat{\sigma}$ of the model parameters when performing experiments with parameter uncertainty using : (a) 500 shots, (b) 1000 shots and (c) $\infty$ shots. The solid and dashed lines represent the convergence in mean error and standard deviation, respectively.
Table 4: Mean and standard deviation (in the brackets) of the parameters using iterative pulse extensions.